OpenStudy (anonymous):

cos __ = 7/17 a) 39 degrees b) 30 degrees c) 26.57 degrees d) 60 degrees

OpenStudy (luigi0210):

Take the inverse cos.. \[\LARGE x=cos^{-1}~(\frac{7}{17})\]

OpenStudy (johnweldon1993):

\[\large \cos \theta = \frac{7}{17}\] \[\large \theta = \cos^{-1}\frac{7}{17}\]

OpenStudy (luigi0210):

^Depends on what your variable was.

OpenStudy (anonymous):

it's actually suppose to be cos __ = 7/14

jigglypuff314 (jigglypuff314):

same idea, just change 17 to 14 :)

OpenStudy (radar):

It is 

OpenStudy (anonymous):

I got the answer 66.6 but its none of the options .

OpenStudy (johnweldon1993):

Are you doing the corrected 7/14 one? I get something different \[\large \theta = \cos^{-1} \frac{7}{14}\]

OpenStudy (anonymous):

yeah , im putting it in too the calculator how you told me to do it , & that's the answering im getting.

OpenStudy (johnweldon1993):

so 7/14 is .5 \[\large \theta = \cos^{-1}(.5)\] I get \[\large \theta = 59.999\]

OpenStudy (anonymous):

oh I was putting 7/14 in , I didn't know I had to make it a decimal first . so the answer would be D ?

OpenStudy (johnweldon1993):

Well yeah you had to do the division first...or at least make sure to put (7/14) But yes D is your answer :)